# Early Jamming Detection in Mobile Indoor Scenarios via Deep Learning

### v1

@@ III System and Adversary Model
Fig. 2 shows the scenario and adversary model considered in this work.
@@
Figure 2: Reference System and Adversary Model. While operating indoors, a drone tries to detect jamming (red color) early, before it might affect the quality of the legitimate communication link (blue color).
@@
We consider an indoor scenario, where drones operate to achieve (semi)-automatized tasks, e.g., warehouse inspection. We consider the existence of a communication link between the drone and a Ground Control Station (GCS), which can be used either to pilot the drone, in the case of a Remotely-Piloted Aircraft System (RPAS), or to report telemetry data, in the case of (semi) autonomous operations. Independently of the usage of the communication link, we do not make any assumption on the nature of the drone operations, which can be either instructed by a human pilot or (semi) autonomous. Without loss of generality, we consider that the communication link between the drone and the GCS adopts the BPSK modulation scheme. This is a reasonable assumption, as such a scheme allows one to mitigate as much as possible the noise affecting indoor communication channels, being used also in modern WiFi standards.
@@
We also consider the deployment of a static jammer in the area, which injects noise into the wireless channel used for communication between the drone and the GCS. We assume that such a jammer continuously emits interfering signals, with the highest possible transmission power, to affect as much as possible ongoing wireless communications in the deployment area. Unlike the contribution in [7], we do not make any assumptions about the specific jamming signal: it can be Additive White Gaussian Noise (AWGN), a single tone, and even deceptive jamming, adopting the same modulation scheme used by the legitimate communication link. Also, being possibly unaware of the sampling rate of the legitimate communication link, the jammer transmits signals with the highest possible sampling ratio, limited only by the hardware used to carry out the attack. Note that when the attacker is unaware of the modulation used by the legitimate communication link, he/she can perform modulation-agnostic jamming, e.g. by injecting AWGN or a single tone, centered on the channel of interest. Instead, if the attacker is aware of the modulation used by the target link, it can use deceptive jamming, i.e., injecting a signal characterized by the same digital modulation (and possibly the same message pattern) of legitimate messages, further complicating the detection process.
@@
Being bounded by the maximum achievable transmission power, the jammer significantly impacts ongoing communications only in a specific area around its location. In fact, the Received Signal Strength (RSS) associated with the jamming signal depends on the distance between the jammer and the receiver location (drone), which is highest in the proximity of the jammer and decreases with further movement. The described wireless propagation effect generates a jammed area, where wireless communication is disrupted. In this context, the mobile receiver (drone), moving toward the jammed area, wants to promptly detect the jamming signal, before the effect of the jamming on the quality of the communication link becomes noticeable, i.e., it causes a significant increase in BER. In fact, such an early jamming detection mechanism would improve drone situational awareness, as it would allow the GCS to be aware of the imminent jamming and take action immediately, without relying on a predefined set of actions (e.g., landing, returning to the starting point).
IV Methodology
@@
This section describes BloodHound$+$, i.e., the methodology we propose to detect jamming before it significantly affects the BER of the legitimate communication link significantly. In summary, BloodHound$+$transforms jamming detection into an anomaly detection problem on images generated by encoding the current state of the legitimate communication channel. Overall, we can identify two main building blocks of our solution: the Image Generation and the Jamming Detection, described below. Tab. I summarizes the main notation used below, with a short description.
TABLE I: Notation and brief description.
@@
Number of samples per image.
@@
Figure 3: Graphical overview of the image generation process of BloodHound$+$. We acquire raw I-Q data through a generic SDR, we plot them through the traditional I-Q plane representation, we compute a bi-variate histogram based on the density of samples in given areas of the plane, and then, we obtain an image.
The input to the process is represented by a sequence of raw I-Q samples. Such samples can be collected using a SDR or any hardware capable of obtaining PHY wireless channel information (e.g., spectrum analyzers). The amount of samples used to generate images, namely $n$, is one of the degrees of freedom of our solution, and it can be configured by trading off the overall accuracy with the computational requirements of the solution (see Section V for a detailed evaluation of the impact of this parameter). We represent the sequence of I-Q samples through the traditional I-Q plane, where we display the component $I$ on the x-axis and the component $Q$ on the y-axis. In a benign scenario (no jamming), such a representation generates a number of clouds of I-Q values, approximately centered on the value of the expected symbol, as explained in Section II-A. Based on such a representation, we build a bivariate histogram. Specifically, we divide the I-Q plane around the cloud of points into tiles $N\times M$, where the values of $N$ and $M$ depend on the dimensions of the images we want to obtain. Then, for each tile $a_{m,n}$, we evaluate the number of I-Q samples that fall into the tile itself. When the number of I-Q samples falling on a tile exceeds the value $255$, we truncate it to the maximum value, to guarantee $a_{m,n}\in[0,255],\forall(m,n)$. We consider the output of such a process as a pixel value. As a consequence, the output of the image generation process is a grayscale image corresponding to the received profile of I-Q samples. Note that, in principle, we can also choose to work with color images (3-D matrices). In Section V we evaluate this configuration, adopted in [8], and show its pros and cons in the context of physical-layer jamming detection.
Jamming Detection. The jamming detection process is the building block of BloodHound$+$ dedicated to the timely detection of possible jamming affecting the wireless communication channel. It is a DL-based process using sparse autoencoders, so involving a training and testing process.
Fig. 4 shows the architecture of the adopted autoencoder.
@@
The input to the autoencoder is represented by the images obtained as a result of applying the image generation process to I-Q samples collected from the wireless channel. At training time, we acquire I-Q samples corresponding to the typical behavior of the communication link. We denote images corresponding to such a scenario configuration as unjammed images. We acquire $n$ I-Q samples and generate from them images of size $M\times N$. In our deployment, $n=10^{5}$ and $M=N=224$, to match the size of images used as input to various NN (we compare our performance to two benchmark solutions using Convolutional (CNN) in Section V). As encoder, we used the logarithmic sigmoid functions, with $K=16$ units (a.k.a. neurons). As a result, we obtain a latent representation of the input image, consisting of $K=16$ dimensions. Such a latent representation summarizes the relevant features of the input image, significantly compressing its dimension (compared to the input image). Then we submit the latent representation vectors to the decoder, using a linear decoder transfer function with a total number of $J=50,176$ neurons. In our context, using the logistic sigmoid activation function in the decoder’s units did not allow our solution to converge to a good solution. We suspect vanishing gradients to be the cause, which is further supported by the better convergence we obtain when using linear activation functions in our decoder layer. In principle, autoencoders allow the use of multiple hidden layers. Overall, the higher the number of hidden layers, the better the performance of the classifier at run-time, but also the higher the computational overhead of the methodology. Here, we use two hidden layers and the sparsity regularization technique, in line with the architecture of sparse nonlinear autoencoders. As we show in Section V, such a choice allows us to obtain remarkable classification accuracy, while achieving a computational cost lower than that of more complex architectures.
This process provides us with a reconstructed image of the same dimension as the input. We first convert the matrix of dimension $m\times n$ in a vector of dimension $d=M\cdot N$, concatenating the rows of the image one after the other. Then, we compute the mse loss function as in Eq. 2 (see Section II-B). During training, we acquire several images corresponding to the regular (expected) behavior of the wireless channel, building a corresponding profile of such a channel when displayed through images. We compute a threshold $\tau$ on such a profile, as explained in Section II-B, to distinguish the regular channel conditions from the unexpected one. At testing time, we compare the mse value obtained from a run-time acquisition of the wireless channel with the threshold $\tau$ previously cited. If the mse of the input image is equal or greater than the threshold, the autoencoder produces a positive prediction, meaning that there is a jammer that affects the communication channel. Otherwise, when mse is lower than the threshold, the autoencoder outputs a negative prediction, meaning there is no jamming. We provide below some more details on the training of the autoencoder and the threshold selection process of BloodHound$+$.
@@
Training the autoencoder. For training the autoencoder, we use only unjammed images, i.e., images generated from I-Q samples acquired when there is no jamming affecting the communication link. We do not use any jammed images, i.e., images obtained from a jammed communication channel. We applied this strategy mainly since jamming can be performed in many different ways, typically unknown and unpredictable to the legitimate parties. Instead, our intuition is that we can build a profile of the expected conditions of the communication link, even in very noisy scenarios, and detect jamming as a deviation from such expected conditions. In this context, to guarantee reliable operations for the autoencoder, it is important to gather I-Q samples that cover the largest possible set of expected conditions of the communication channel. In fact, the reliability of the auto-encoder in identifying anomalies leading to jamming depends on the variety of conditions affecting the communication channel, thus reducing false positive events.
@@
Threshold selection. Optimal selection of the decision threshold of an autoencoder is an actively researched problem, which does not have a universally optimal solution [14, 13]. In this work, we adopt the approach suggested by the authors in [12], i.e., we compute the threshold according to Eq. 3.
@@
being ${MSE}_{{train}}$ the set of mses that the autoencoder compute on the training data, ${mean}$ the statistical average and ${std}$ the standard deviation. As demonstrated by the authors in [12] and confirmed by the authors in [13], such a choice is reasonable in scenarios where no anomalous samples are used in the training phase, as in our scenario. Moreover, as acknowledged by the authors in [13], this choice allows us to reduce false negatives when compared to the common choice, i.e., setting $\tau$ to the maximum mse observed on the training samples. In turn, such a choice allows us to boost the chances of detecting jamming earlier. Figure 5 reports an example of the threshold selection process on real data acquired with jammer jamming with $RJP=0.6$ at a distance of $10$ meters from the receiver (see Section V-A for details). We notice that the distribution of the mse values for unjammed images is characterized by smaller values compared to the one of jammed images, with only minimal overlapping on the tails of the distributions. Setting the threshold according to Eq. 3 allows one to reduce false negatives, without affecting performance.
@@
Finally, note that deploying an autoencoder for image-based jamming detection involves setting optimal values of the hyperparameters of such a tool. We discuss the selection of auto-encoder hyperparameters in Section V.
V Experimental Assessment
In this section, we provide the details of our extensive experimental assessment, carried out to evaluate BloodHound$+$ in real indoor scenarios. We introduce the actual measurements used for the following analysis in Section V-A, while in Section V-B we describe the experimental settings. Then Section V-C reports the performance of our approach and compares it with two solutions available in the literature. We extend such results further, evaluating BloodHound$+$ with different hardware and various channel sampling rates (Section V-D), and investigating the capability of identifying deceptive jamming (Section V-E).
V-A Measurements
@@
All of the measurements discussed below have been acquired in an indoor office environment during a regular working day, with people moving around and possibly across the legitimate communication link. Such conditions match, as much as possible, those described in Section III.
@@
The general setup of our measurements includes three entities, i.e., a transmitter, a receiver, and a jammer. In the first set of experiments, we considered SDR Ettus Research X310 [16] featuring a daughterboard UBX160 as reference hardware for all the entities involved in our measurement campaign. Here, we placed the transmitter and the jammer close to each other, while moving the receiver at various distances from the transmitter (see below). For the second set of experiments, we used different hardware, that is the LimeSDR [17]. It is a low-cost, open-source, SDR platform that can be used to support any wireless communication standard. For these experiments, we placed the transmitter and receiver $3$ meters away from each other and placed the jammer between them, at a distance of $1.5$ meters from both entities. We connected the SDRs either via Ethernet (Ettus X310) or USB 2.0 (LimeSDR) to two laptops, one controlling the data transmission and jamming processes and the other taking care of data reception. The received I-Q samples were stored on the laptop connected to the receiving SDR, and subsequently, uploaded to a centralized server for data analysis. Specifically, we used the High Processing Cluster (HPC) available at TU/e, Eindhoven, The Netherlands, providing a CPU E2124 with 4 cores running at 3.3 GHz and 32 GB of RAM, as well as 2 GPUs Tesla V1000 running at 32 with 256 GB of RAM.
@@
Regarding the software, for both setups, we used the GNURadio v3.8 development toolkit [18]. We set the carrier frequency $f_{c}=900$ MHz for both the legitimate communication link and the jamming. The transmitter and receiver are configured to exchange packets containing a repeating sequence of 256 bytes, encoded by a Constellation Modulation block using the regular BPSK modulation scheme.
@@
For the first setup using the SDR Ettus X310, we configured a sample rate of 1M samples per second at the transmitter, the receiver, and the jammer. We set the normalized transmission power and receiver gain to the maximum value of $1$, corresponding to approximately 15 dBm (32 mW) of transmission power. At the receiver, we set up the reception chain of the BPSK modulation, including (i) an Adaptive Gain Control (AGC) block, to mitigate the signal level fluctuations introduced by the multipath fading; (ii) a Symbol Sync, which performs timing synchronization; (iii) a Costas Loop, which locks to the center frequency of the signal and down-converts it to baseband; and, (iv) finally, a Constellation Decoder block, which decodes the constellation points. We saved the I-Q data obtained as the output of the Constellation Decoder block. We did not use any channel estimation techniques, in order not to filter out any channel effects useful for jamming detection. Regarding the jammer, we chained two blocks: (i) a signal source, which can be an analog $\sin$ wave (tone jammer) or a digital sequence of Gaussian-distributed values (Gaussian jammer); and (ii) the USRP Sink block, which sends the signal to the radio for actual transmission. To emulate the scenario described in Section III, we placed the entities at different distances and, to further emulate the movement, we changed the relative values of the jammer transmission power between 00 and $0.8$, i.e., between 0 and 7.94mW (9dBm), respectively. Values greater than $0.8$ cause a complete disruption of the BER of the signal (see Section V-C), making our solution unnecessary. In fact, when using a static setup, the reduction of the transmission power of the jammer makes the received jamming power level at the receiver weaker, allowing us to investigate the effect of a jammer located further away from the communication link.
@@
For the second setup using the LimeSDR, we considered different values of sample rate at the receiver and at the jammer. In particular, consider the formula $t_{s}=K\times t_{S,R}$, where $t_{S,R}$ is the reference sample rate of 1M samples per second, $t_{s}$ is the actual sample rate used in the measurement, and $K$ an oversampling factor. We carried out experiments considering different values of the oversampling ratio both at the receiver and at the jammer, namely, the Receiver Oversampling Ratio (ROR) and Jamming Oversampling Ratio (JOR). Specifically, we varied both the ROR and the JOR in the range $[1,4]$. For all such experiments, we tested two jamming strategies: tone jamming, i.e., jamming with a sinusoid signal, and deceptive jamming, i.e., jamming with exactly the same bit-sequence BPSK-modulated signal delivered between the legitimate transmitter and the receiver.
@@
Overall, the two measurement setups described above allowed us to investigate the effectiveness of BloodHound$+$ while varying a large range of configuration parameters including: (i) the RJP at the receiver, (ii) the distance of the jammer from the legitimate communication link, (iii) the oversampling ratios at the receiver and the jammer, and (iv) the types of jamming and the type of radios used for the experiments.
V-B Experimental Settings

### v4

@@ IV System and Adversary Model
Fig. 2 shows the scenario and adversary model considered in this work, inspired by the real-world use case discussed in [9].
@@
Figure 2: Reference Scenario. While operating indoors, a drone tries to detect jamming (red color) in a low-BER regime, i.e., before the jamming affects the quality of the legitimate communication link (blue color).
@@
We consider an indoor scenario where drones operate to achieve (semi)-automatized tasks, e.g., warehouse inspection or home surveillance. The relevance of such a scenario is confirmed by recent news, reporting that leading companies such as IKEA and Amazon are already experimenting with commercial products for this purpose [9][10]. We consider the existence of a communication link between the drone and a Ground Control Station (GCS), which can be used either to pilot the drone, in the case of a Remotely-Piloted Aircraft System (RPAS), or to report telemetry data, in the case of (semi) autonomous operations [46]. Independently of the communication link usage, we do not make any assumption on the nature of the drone operations, which can be either instructed by a human pilot or (semi-) autonomous. Also, we do not make assumptions on the presence of a positioning technology, besides that it is not affected by spoofing [47]. Without loss of generality, we consider that the communication link between the drone and the GCS adopts the BPSK modulation scheme. The mentioned assumption is reasonable, as such a scheme allows one to mitigate the noise affecting indoor communication channels as much as possible, being also used in modern WiFi standards.
@@
We also consider the deployment of a static jammer in the area, which injects noise into the wireless channel used for communication between the drone and the GCS. We assume that such a jammer continuously emits interfering signals with the highest possible transmission power, to affect ongoing wireless communications in the deployment area as much as possible. Unlike the contribution in [16], we do not make any assumptions about the specific jamming signal: it can be Additive White Gaussian Noise (AWGN), a single tone, and even deceptive jamming, adopting the same modulation scheme used by the legitimate communication link. Also, being possibly unaware of the sampling rate of the legitimate communication link, the jammer transmits signals with the highest possible sampling ratio, limited only by the hardware used to carry out the attack. Note that when the attacker is unaware of the modulation used by the legitimate communication link, they can perform modulation-agnostic jamming, e.g. by injecting AWGN or a single tone centered on the channel of interest. Instead, suppose that the attacker is aware of the modulation used by the target link. In that case, it can use deceptive jamming, i.e., injecting a signal characterized by the same digital modulation (and possibly the same message pattern) of legitimate messages, further complicating the detection process.
@@
Being bounded by the maximum achievable transmission power, the jammer significantly impacts ongoing communications only in a specific area around its location. In fact, the RSS associated with the jamming signal depends on the distance between the jammer and the receiver location (drone), which is highest in the proximity of the jammer and decreases with further movement. The described wireless propagation effect generates a jammed area, disrupting wireless communications. In this context, the mobile receiver (drone), moving toward the jammed area, wants to promptly detect the jamming signal in a low-BER regime, i.e., before the effect of the jamming on the quality of the communication link becomes noticeable, causing a significant increase in the BER. In fact, such a jamming detection mechanism would improve drone situational awareness, as it would allow GCS to be aware of imminent jamming and take action immediately without relying on a predefined set of steps (e.g., landing, returning to the starting point).
V Methodology
@@
This section describes BloodHound$+$, i.e., the methodology we propose to detect jamming in a low-BER regime. In summary, BloodHound$+$transforms jamming detection into an anomaly detection problem on images generated by encoding the current state of the communication channel. Overall, we can identify two main building blocks of our solution: the Image Generation and the Jamming Detection, described below. Tab. II summarizes the primary notation used below, with a short description.
TABLE II: Notation and brief description.
@@
Number of I-Q samples per image.
@@
Figure 3: Graphical overview of the image generation process of BloodHound$+$. We acquire raw I-Q data through a generic SDR, plot them through the traditional I-Q plane representation, compute a bi-variate histogram based on the density of samples in given areas of the plane, and then obtain an image.
The input to the process is represented by a sequence of raw I-Q samples. Such samples can be collected using a SDR or any hardware capable of obtaining PHY wireless channel information (e.g., spectrum analyzers). The amount of samples used to generate images, namely $n$, is one of the degrees of freedom of our solution and can be configured by trading off the overall accuracy with the computational requirements of the solution (see Section VI for a detailed evaluation of the impact of this parameter). We represent the sequence of I-Q samples through the traditional I-Q plane, where we display the component $I$ on the x-axis and the component $Q$ on the y-axis. In a benign scenario (no jamming), such a representation generates several clouds of I-Q values, approximately centred on the value of the expected symbol, as explained in Section III-A. Based on this representation, we build a bi-variate histogram. Specifically, we divide the I-Q plane around the cloud of points into tiles $N\times M$, where the values of $N$ and $M$ depend on the dimensions of the images we want to obtain. Then, for each tile $a_{m,n}$, we evaluate the number of I-Q samples that fall into the tile itself. When the number of I-Q samples falling on a tile exceeds the value $255$, we truncate it to the maximum value, to ensure $a_{m,n}\in[0,255],\forall(m,n)$. We consider the output of such a process as a pixel value. As a consequence, the output of the image generation process is a grayscale image corresponding to the received profile of I-Q samples. Note that, in principle, we might also work with colored images (3-D matrices). In Section VI, we evaluate this configuration, adopted in [17], and show its pros and cons in physical-layer jamming detection.
Jamming Detection. The jamming detection process is the building block of BloodHound$+$ dedicated to the timely detection of possible jamming affecting the wireless communication channel. It is a DL-based process using sparse autoencoders, so it involves a training and testing process. We highlight that our manuscript does not aim to provide a new autoencoder architecture. Instead, as discussed in the related work (Sect. II) and summarized in Tab. I, the innovation of our manuscript is that BloodHound$+$ is the first solution applying autoencoders to solve a jamming detection problem. Fig. 4 shows the architecture of the adopted autoencoder.
@@
The input to the autoencoder is represented by the images obtained as a result of applying the image generation process to I-Q samples collected from the wireless channel. At training time, we acquire I-Q samples corresponding to the typical behavior of the communication link. We denote images corresponding to such a scenario configuration as unjammed images. We acquire $n$ I-Q samples and generate images of size $M\times N$. In our deployment, $n=10^{5}$ and $M=N=224$, to match the size of images used as input to various NN (we compare our performance to two benchmark solutions using CNN in Section VI). As encoder, we used logarithmic sigmoid functions with $K=16$ units (a.k.a. neurons). As a result, we obtain a latent representation of the input image consisting of $K=16$ dimensions. Such a latent representation summarizes the relevant features of the input image, significantly compressing its dimension (compared to the input image). Then, we submit the latent representation vectors to the decoder using a linear decoder transfer function with a total number of $J=50,176$ neurons. In our context, using the logistic sigmoid activation function in the decoder’s units did not allow our solution to converge to a good solution. We suspect vanishing gradients to be the cause, further supported by the better convergence we obtain when using linear activation functions in our decoder layer. In principle, autoencoders allow the use of multiple hidden layers. Overall, the higher the number of hidden layers, the better the performance of the classifier at run-time, but also the higher the computational overhead of the methodology. Here, we use two hidden layers and the sparsity regularization technique, in line with the architecture of sparse nonlinear autoencoders. As we show in Section VI, such a choice allows us to obtain remarkable classification accuracy while achieving a computational cost lower than that of more complex architectures. This process provides a reconstructed image of the same dimension as the input. We first convert the matrix of size $m\times n$ into a vector of dimension $d=M\cdot N$, concatenating the rows of the image one after the other. Then, we compute the mse loss function as in Eq. 2 (see Section III-B). During training, we acquire several images corresponding to the regular (expected) behavior of the wireless channel, building a corresponding profile of such a channel when displayed through images (i.e., our hypothesis). We compute a threshold $\tau$ on such a profile, as explained in Section III-B, to distinguish the regular channel conditions from the unexpected one. At testing time, we compare the mse value obtained from a run-time acquisition of the wireless channel with the threshold $\tau$ previously cited. Suppose that the mse of the input image is equal to or greater than the threshold. In that case, the autoencoder produces a positive prediction, meaning that a jammer affects the communication channel. Otherwise, when mse is lower than the threshold, the autoencoder outputs a negative prediction, meaning there is no jamming. In line with the logic of any ML and DL solution, the rationale of BloodHound$+$ aligns with statistical hypothesis testing methods. In the following, we provide more details on the autoencoder training process and the threshold selection methodology.
@@
Training the autoencoder. For training the autoencoder, we use only unjammed images, i.e., images generated from I-Q samples acquired when no jamming affects the communication link. We do not use any jammed images, i.e., images obtained from a jammed communication channel. We applied this strategy mainly since jamming can be performed in many different ways, typically unknown and unpredictable to the legitimate parties. Instead, our intuition is that we can build a profile of the expected conditions of the communication link, even in very noisy scenarios, and detect jamming as a deviation from such expected conditions. In this context, to guarantee reliable operations for the autoencoder, it is crucial to gather I-Q samples that cover the most extensive possible set of expected conditions of the communication channel. In fact, the reliability of the auto-encoder in identifying anomalies leading to jamming depends on the variety of conditions affecting the communication channel, thus reducing false positive events.
@@
Threshold selection. Optimal selection of the decision threshold of an autoencoder is an actively researched problem, which has yet to have a universally optimal solution [48, 45]. In this work, we adopt the approach suggested by the authors in [44], i.e., we compute the threshold according to Eq. 3.
@@
being ${MSE}_{{train}}$ the set of mses that the autoencoder compute on the training data, ${mean}$ the statistical average and ${std}$ the standard deviation. As demonstrated by the authors in [44] and confirmed by the authors in [45], such a choice is reasonable in scenarios where no anomalous samples are used in the training phase, as in our scenario. Furthermore, as acknowledged by the authors in [45], this choice reduces false negatives compared to the standard option, i.e., setting $\tau$ to the maximum mse observed in the training samples. In turn, such a choice increases the chances of detecting jamming in a low-BER regime. Figure 5 reports an example of the threshold selection process on actual data acquired with jammer jamming with $RJP=0.6$ at a distance of $10$ meters from the receiver (see Section VI-A for details). We notice that the distribution of the mse values for unjammed images is characterized by smaller values compared to the one of jammed images, with only minimal overlap at the tails of the distributions. Setting the threshold according to Eq. 3 allows one to reduce false negatives without affecting performance.
@@
Finally, note that deploying an autoencoder for image-based jamming detection involves setting optimal values of the hyperparameters of such a tool. We discuss the selection of auto-encoder hyperparameters in Section VI.
VI Experimental Assessment
In this section, we provide the details of our extensive experimental assessment, carried out to evaluate BloodHound$+$ in real indoor scenarios. We introduce the actual measurements used for the following analysis in Section VI-A, while in Section VI-B we describe the experimental settings. Then Section VI-C reports the performance of our approach and compares it with two solutions available in the literature. We extend such results further, evaluating BloodHound$+$ with different hardware and various channel sampling rates (Section VI-D) and investigating the capability of identifying deceptive jamming (Section VI-E).
VI-A Measurements
@@
All the measurements discussed below have been acquired in an indoor office environment during a regular working day, with people moving around and possibly across the legitimate communication link. Such conditions match, as much as possible, those described in Section IV.
@@
The general setup of our measurements includes three entities, i.e., a transmitter, a receiver, and a jammer. In the first set of experiments, we considered SDR Ettus Research X310 [50] featuring a daughterboard UBX160 as reference hardware for all the entities involved in our measurement campaign. Here, we placed the transmitter and the jammer close to each other while moving the receiver at various distances from the transmitter (see below). For the second set of experiments, we used different hardware, i.e., the LimeSDR [51]. It is a low-cost, open-source, SDR platform supporting any wireless communication standard. For these experiments, we placed the transmitter and receiver $3$ meters away from each other and placed the jammer between them at a distance of $1.5$ meters from both entities. We connected the SDRs either via Ethernet (Ettus X310) or USB 2.0 (LimeSDR) to two laptops, one controlling the data transmission and jamming processes and the other taking care of data reception. The received I-Q samples were stored on the laptop connected to the receiving SDR and subsequently uploaded to a centralized server for data analysis. Specifically, we used the High-Performance Computing (HPC) cluster available at TU/e, Eindhoven, The Netherlands, providing a CPU E2124 with four cores running at 3.3 GHz and 32 GB of RAM, as well as 2 GPUs Tesla V1000 running at 32 with 256 GB of RAM.
@@
Regarding software, for both setups, we used the GNURadio v3.8 development toolkit [52]. We set the carrier frequency $f_{c}=900$ MHz for both the legitimate communication link and the jamming. We configured the transmitter and receiver to exchange packets containing a repeating sequence of 256 bytes, encoded by a Constellation Modulation block using the regular BPSK modulation scheme.
@@
For the first setup using the SDR Ettus X310, we configured a sample rate of 1M samples per second at the transmitter, the receiver, and the jammer. We set the normalized transmission power and receiver gain to the maximum value of $1$, corresponding to approximately 15 dBm (32 mW) of transmission power. At the receiver, we set up the reception chain of the BPSK modulation, including (i) an Adaptive Gain Control (AGC) block, to mitigate the signal level fluctuations introduced by the multipath fading; (ii) a Symbol Sync, which performs timing synchronization; (iii) a Costas Loop, which locks to the center frequency of the signal and down-converts it to baseband; and, (iv) finally, a Constellation Decoder block, which decodes the constellation points. We saved the I-Q data obtained as the output of the Constellation Decoder block. We did not use any channel estimation techniques to filter out any channel effects beneficial for jamming detection. Regarding the jammer, we chained two blocks: (i) a signal source, which can be an analog $\sin$ wave (tone jammer) or a digital sequence of Gaussian-distributed values (Gaussian jammer); and (ii) the USRP Sink block, which sends the signal to the radio for actual transmission. To emulate the scenario described in Section IV, we placed the entities at different distances and, to further mimic the movement, we changed the relative values of the jammer transmission power between 00 and $0.8$, i.e., between 0 and 7.94mW (9dBm), respectively. Values greater than $0.8$ cause a complete disruption of the BER of the signal (see Section VI-C), making our solution unnecessary. In fact, when using a static setup, the reduction of the transmission power of the jammer makes the received jamming power level at the receiver weaker, allowing us to investigate the effect of a jammer located further away from the communication link.
@@
For the second setup using the LimeSDR, we considered different sample rate values at the receiver and the jammer. In particular, consider the formula $t_{s}=K\times t_{S,R}$, where $t_{S,R}$ is the reference sample rate of 1M samples per second, $t_{s}$ is the actual sample rate used in the measurement, and $K$ an oversampling factor. We carried out experiments considering different values of the oversampling ratio both at the receiver and at the jammer, namely, the Receiver Oversampling Ratio (ROR) and Jamming Oversampling Ratio (JOR). Specifically, we varied both the ROR and the JOR in the range $[1,4]$. For all such experiments, we tested two jamming strategies: tone jamming, i.e., jamming with a sinusoid signal, and deceptive jamming, i.e., jamming with precisely the same bit-sequence BPSK-modulated signal delivered between the legitimate transmitter and the receiver.
@@
In general, the two measurement setups described above allowed us to investigate the effectiveness of BloodHound$+$ while varying an extensive range of configuration parameters, including: (i) the RJP at the receiver, (ii) the distance of the jammer from the legitimate communication link, (iii) the oversampling ratios at the receiver and the jammer, and (iv) the types of jamming and the type of radios used for the experiments.
VI-B Experimental Settings